An athelete takes 2s to reach his maximum speed of 18km/hr . What is the magnitude  of his average accelaretion

Grade:6

1 Answers

Saransh Anand
7 years ago
v = 18km/hr = 5m/s 
u=0 m/s
t = 2 s 
a = ?
 
According to equation of motion 
v = u + a.t 
5 = 0 + a.2 
a = 5/2
   = 2.5 m/s2